OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (anonymous):

y - 9 = 3(x - 4)

OpenStudy (anonymous):

the equation is y- y1 = m (x - x1)

OpenStudy (anonymous):